# High-dimensional sliced inverse regression with endogeneity

### v1

@@ 6.1 Omitted variables: Mouse obesity data
We applied the proposed 2SLSIR estimator to perform joint SDR and variable selection on a mouse obesity dataset from Wang et al., (2006), and compared the results obtained to those from Lin et al., (2015) who analyzed the same data using their two-stage Lasso (2SLasso) estimator, along with the single stage Lasso SIR (LSIR) estimator. Briefly, the study comprised 334 mice fed a high-fat diet, and who were then genotyped using 1327 single nucleotide polymorphisms (SNPs) with gene expressions profiled on microarrays that include probes for $23,388$ genes. We followed the same pre-processing steps as in Lin et al., (2015), and obtained a complete dataset with $p=2825$ genes and $q=1250$ SNPs on $n=287$ mice. We treated the SNPs as instruments ($\mathbf{Z}$), and focused on identifying important genes ($\mathbf{X}$) related to the body weight ($y$). As mentioned in Section 1, the relationship between gene expression covariates and the outcome is likely confounded by omitted variables, experimental conditions, and environmental perturbations. As such, endogeneity due to omitted variable bias is expected to be a critical issue in this application. As justified in Lin et al., (2015), genetic variants (represented by SNP) are valid instruments because the mice in this study come from the same genetic cross.
@@
To select the number of structural dimensions $d$, we used $k$-means clustering of the (normalized) eigenvalues of the estimated matrix $\widehat{\bm{\Lambda}}$ similar to that of Lin et al., (2019), with 2SLSIR and 2SLasso both selecting $d=1$. Next, for each of the three estimators, we applied stability selection to compute the empirical selection probability (Meinshausen and Bühlmann,, 2010) for each gene over 100 random subsamples of size $\lfloor n/2\rfloor$ at each value of the tuning parameter. The upper bound for the per-family error rate was one while the selection cutoff was set to $0.75$, meaning up to 75% number of components can be included in the estimates in each subsample. The results in Figure 1 present the selection path for each gene as a function of the second-stage tuning parameter $\mu_{2}$. With the maximum selection probability across all the tuning parameters set to $0.5$, LSIR selected seven genes, 2SLasso selected nine genes, while the proposed 2SLSIR selected six genes. Only one gene was chosen with a very high selection probability across all the three methods, while one additional gene is commonly selected by 2SLSIR and 2SLasso.
@@
Figure 1: Stability selection paths for three different methods applied to the mouse obesity data based on 100 random subsamples. With the maximum selection probability of at least 0.5, the gene selected by all the three methods is displayed in red, while the additional gene only selected by the two-stage methods is displayed in blue.
To further interpret the biological consequences of the selected genes, we conducted a Gene Ontology (GO) enrichment analysis for the genes chosen by each method. By annotating these genes for their participation in gene ontology, this analysis aims to identify which biological functions, known as GO terms, are over-represented in the selected genes (Bleazard et al.,, 2015). This over-representation is measured by the $p$-values, adjusted for multiple comparisons, of the null hypothesis that this gene set is randomly associated with the GO, so a smaller (larger) $p$-value indicates a more significant biological function. Furthermore, the gene ontology is divided into three main groups: biological processes (BP), molecular function (MF), and cellular component (CC). Figure 2 indicates that all the three methods capture a strong association between the obesity and lipid-related processes, including lipid transport, metabolism, and regulation. The one-stage LSIR focuses on genes regulating lipid transport and metabolic processes (e.g., lipid transporter activity, monocarboxylic acid transport), linking BMI to lipid dynamics at the cellular level. By contrast, both the 2SLasso and 2SLSIR estimators emphasize plasma lipoprotein metabolism (e.g., plasma lipoprotein particle clearance, protein-lipid complex). However, while the former puts more emphasis on the interplay between lipid metabolism and immune regulation (e.g., immunoglobulin binding), the latter focuses more on lipid catabolism and extracellular lipid transport (e.g., extracellular matrix structural constituent, lipase activity).
@@ 6.2 Measurement error: NHANES data
The National Health and Nutrition Examination Survey (NHANES) survey aims to assess the health and nutritional status of individuals in the United States, tracking the evolution of this over time (Centers for Disease Control and Prevention,, 2018). During the 2009-2010 survey period, participants were interviewed and asked to provide their demographic background, nutritional habits, and to undertake a series of health examinations. To assess the nutritional habits of participants, dietary data were collected using two 24-hour recall interviews, wherein the participants self-reported the consumed amounts of a set of food items during the 24 hours before each interview. Based on these recalls, daily aggregated consumption of water, food energy, and other nutrition components such as total fat and total sugar consumption were computed. In this section, we revisit the data application in Nghiem et al., 2023a , seeking to understand the relationship between total cholesterol level ($y$) for $n=3343$ women and their long-term intakes of $p=42$ nutrition components, particularly a linear combination/s of these components carrying the full regression information, and whether these combinations are sparse in nature.
@@
Let $\mathbf{W}_{ij}$ denote the vector of 42 nutritional covariates at the $j$th interview for $j=1,2$. Dietary recall data are well-known to be subject to measurement errors, so we assume $\mathbf{W}_{ij}=\mathbf{X}_{i}+\mathbf{E}_{ij}$, where $\mathbf{X}_{i}$ represents the $i$th latent long-term nutritional intake, and $\mathbf{E}_{ij}$ denotes an independent additive measurement error term following a multivariate Gaussian distribution with zero mean and covariance $\bm{\Sigma}_{u}$. We used an instrumental variable approach to account for this measurement error. Specifically, following Carroll et al., (Section 6.5.1, 2006) we treated the first replicate $\mathbf{W}_{i1}=\mathbf{W}_{i}$ as a surrogate for $\mathbf{X}_{i}$ and the second replicate $\mathbf{W}_{i2}=\mathbf{Z}_{i}$ as an instrument. Note in this case, the number of instruments $q=p$. We then applied the linear instrumental variable model $\mathbf{X}_{i}=\bm{\Gamma}^{\top}\mathbf{Z}_{i}+\mathbf{U}_{i}$ where $\bm{\Gamma}$ is allowed to be a $q\times p$ column-sparse matrix. It follows that we can write $\mathbf{W}_{i}=\bm{\Gamma}^{\top}\mathbf{Z}_{i}+\tilde{\mathbf{U}}_{i}$ where $\tilde{\mathbf{U}}_{i}=\mathbf{U}_{i}+\mathbf{E}_{i1}$. We then applied the proposed 2SLSIR estimator with the true covariate matrix $\mathcal{X}$ replaced by the $n\times p$ matrix $\mathcal{W}$, where we selected $d=1$ as the structural dimension following Nghiem et al., 2023a .
We first examined the estimated matrix coefficients $\widehat{\bm{\Gamma}}$ from the first stage of the 2SLSIR estimation procedure. Figure S1 in the Supplementary Materials shows that the biggest element in each column of $\widehat{\bm{\Gamma}}$ is typically the diagonal element, so the most important instrument to model the consumption of any dietary component in the second interview is, unsurprisingly, the consumption of the same component in the first interview. Overall, each column of $\widehat{\bm{\Gamma}}$ is relatively sparse, with the number of non-zero coefficients in each column ranging from 3 to 20.
@@
Next, we applied the same stability selection approach as in Section 6.1 to the 2SLSIR estimator and compared the resulting paths with the LSIR and 2SLasso estimators. With the maximum selection probability across all the tuning parameters being at least 0.5, LSIR selected five nutritional covariates, 2SLasso selected five covariates, while the proposed 2SLSIR selected six covariates. Figure 3 presents the selection path as a function of the tuning parameter $\mu_{2}$ for each nutrition component. The two components commonly selected by all the three methods were vitamin D and vitamin K, while both 2SLSIR and 2SLasso additionally select one more common component in folic acid. Vitamin D is the component with the largest estimated coefficients in both the 2SLSIR and 2SLasso estimators, and was the second largest in the LSIR estimator.
@@ 7 Discussion
In this paper, we propose an innovative two-stage Lasso SIR method to perform simultaneous sufficient dimension reduction and variable selection when endogeneity is present. We establish theoretical bounds for estimation and selection consistency of the proposed estimator, allowing the dimensionality of both the covariates and instruments to grow exponentially with sample size. Our simulation studies demonstrate that the 2SLSIR estimator exhibits superior performance compared with the one-stage Lasso SIR method that ignores endogeneity, and the two-stage Lasso method, that ignores non-linearity in the outcome model. We apply the proposed method in two data applications where endogeneity is well-known to be present, one in nutrition and one in statistical genomics.
Future research can extend the proposed two-stage Lasso SIR method to accommodate a possible non-linear relationship between the covariates and the instruments in the first stage. Different sparsity structures on the coefficient matrix $\bm{\Gamma}$, such as a low-rank matrix, can also be considered. One can also examine endogeneity in other sufficient dimension reduction methods, such as inverse regression models Cook and Forzani, (2008); Bura et al., (2022). Finally, the methods can be adapted to more complex settings, such as clustered or longitudinal data, for example as in (Nghiem and Hui,, 2024).
Acknowledgments
This work was supported by the Australian Research Council Discovery Project Grant DP230101908. Thanks to Shila Ghazanfar for useful discussions regarding the mouse obesity data application.
Supplementary Materials
S1 Performance of Lasso SIR under endogenity
@@ Acknowledgments
$\begin{bmatrix}\mathbf{X}\\ \varepsilon\\ \end{bmatrix}\sim N_{p+1}\left(\bm{0},\begin{bmatrix}\mathbf{I}_{p}&\bm{\Sigma}_{\mathbf{X},\varepsilon}\\ \bm{\Sigma}_{\mathbf{X},\varepsilon}^{\top}&1\end{bmatrix}\right),$ (12)
@@
$\begin{bmatrix}\mathbf{X}\\ y\\ \end{bmatrix}\sim N_{p+1}\left(\bm{0},\begin{bmatrix}\mathbf{I}_{p}&\bm{\beta}+\bm{\Sigma}_{\mathbf{X},\varepsilon}\\ \bm{\beta}^{\top}+\bm{\Sigma}_{\mathbf{X},\varepsilon}^{\top}&\bm{\beta}^{\top}\bm{\beta}+2\bm{\Sigma}_{\mathbf{X},\varepsilon}\bm{\beta}+1\end{bmatrix}\right),$
@@
For each generated sample, we computed the Lasso SIR estimator $\hat{\bm{\beta}}^{*}$ with the tuning parameter $\mu$ selected by 10-fold cross-validation, as implemented by the LassoSIR package. We reported the estimation error in estimating the column space spanned by $\bm{\beta}$, as measured by $\left\lVert\mathcal{P}(\bm{\beta})-\mathcal{P}(\hat{\bm{\beta}})\right\rVert_{F}$. For variable selection performance, we computed the area under the receiver operating curve (AUC); an AUC of one indicates a clear separation between the estimates of non-important variables and the those of important variables in $\hat{\bm{\beta}}^{*}$.
Table S3 shows that the performance of the Lasso SIR estimator aligns with our theoretical results. In the first case, the correlation vector $\bm{\Sigma}_{\mathbf{X},\varepsilon}$ is proportional to the true $\bm{\beta}$, so the Lasso SIR maintains both estimation and variable selection consistency, as evidenced by a decreasing estimation error when $n$ increases and an AUC of 1. In the second case, the Lasso SIR estimator gives a consistent estimate of the space spanned by $\bm{\eta}^{*}=(1.5,0.5,0,0)/\sqrt{2.25}$, which has a different direction but the same support compared to the true $\bm{\beta}$. Hence, the Lasso SIR is not estimation consistent, but still has an AUC of 1. In the last case, the Lasso SIR estimator gives a consistent estimate for the space spanned by $\bm{\eta}^{*}=(0.5,0.5,0.5,0.5)^{\top}$, which has both a different direction and a different support from those of the true $\bm{\beta}$. As a result, the Lasso SIR estimator is not selection or estimation consistent.
Table S3: Performance of the Lasso SIR estimator under endogeneity, as measured by the estimation error $\left\lVert\mathcal{P}(\bm{\beta})-\mathcal{P}(\hat{\bm{\beta}}^{*})\right\rVert_{F}$ and variable selection performance (AUC).
@@
S2 Proofs
S2.1 Proof of Proposition 1
Without loss of generality, assume the expectation of $\text{E}(\mathbf{X},\varepsilon)=\bm{0}$, the marginal covariance matrix $\bm{\Sigma}_{X}=\mathbf{I}_{p}$, and the marginal variance $\operatorname{\text{Var}}(\varepsilon)=1$. In this case, it follows that
$\begin{bmatrix}\mathbf{X}\\ \varepsilon\\ \end{bmatrix}\sim N_{p+1}\left(\bm{0},\begin{bmatrix}\mathbf{I}_{p}&\bm{\Sigma}_{\mathbf{X},\varepsilon}\\ \bm{\Sigma}_{\mathbf{X},\varepsilon}^{\top}&1\end{bmatrix}\right).$
@@
$\frac{1}{n}\left\lVert\tilde{\bm{y}}^{*}_{k}-\mathcal{X}\tilde{\bm{\beta}}^{*}_{k}-\mathcal{X}\bm{\delta}_{k}^{*}\right\rVert^{2}+\mu_{k}\|\tilde{\bm{\beta}}^{*}_{k}+\bm{\delta}_{k}^{*}\|_{1}\leq\frac{1}{n}\left\lVert\tilde{\bm{y}}^{*}_{k}-\mathcal{X}{\bm{\beta}}^{*}_{k}\right\rVert^{2}+\mu_{k}\|\tilde{\bm{\beta}}^{*}_{k}\|_{1}.$
@@
[…]
@@
[…]
S2.3 Proof of Lemma 1
The instrument $\mathbf{Z}$ is independent of both $\varepsilon$ and $\mathbf{U}$, so the conditional expectation $\hat{\mathbf{X}}=E(\mathbf{X}\mid\mathbf{Z})=\bm{\Gamma}\mathbf{Z}$ is also independent of both $\varepsilon$ and $\mathbf{U}$. By the properties of conditional expectation, we have

### v2

@@ 7.1 Omitted variables: Mouse obesity data
We applied the proposed two-stage lasso SIR estimator to perform simultaneous SDR and variable selection on a mouse obesity dataset from Wang et al., (2006), and compared the results obtained to those from Lin et al., (2015) who analyzed the same data using their two-stage lasso estimator, as well as results based on applying the lasso SIR estimator. Briefly, the study comprised $334$ mice fed a high-fat diet, and who were then genotyped using $1327$ single nucleotide polymorphisms (SNPs) with gene expressions profiled on microarrays that include probes for $23,388$ genes. We followed the same pre-processing steps as in Lin et al., (2015), and obtained a complete dataset with $p=2825$ genes and $q=1250$ SNPs on $n=287$ mice. We treated the SNPs as instruments ($\mathbf{Z}$), and focused on identifying important genes ($\mathbf{X}$) related to the body weight ($y$). As mentioned in Section 1, the relationship between gene expression covariates and the outcome is likely confounded by omitted variables, experimental conditions, and environmental perturbations. As such, endogeneity due to omitted variable bias is expected to be a critical issue in this application. Lin et al., (2015) argued that genetic variants as represented by SNPs are valid instruments because the mice in this study come from the same genetic cross.
@@
To select the number of structural dimensions $d$, we applied the three approaches described in Section 6, and all selected $\hat{d}=1$. Next, for each of the two-stage lasso, one-stage lasso SIR, and two-stage lasso SIR estimators, we applied stability selection to compute the empirical selection probability (Meinshausen and Bühlmann,, 2010) for each gene over 100 random subsamples of size $\lfloor n/2\rfloor$ at each value of the tuning parameter. The upper bound for the per-family error rate was set to one, while the selection cutoff was set to $0.75$, meaning up to 75% number of components can be included in the estimates in each subsample. The results in Figure 1 present the selection path for each gene as a function of the second-stage tuning parameter $\mu_{2}$. With the maximum selection probability across all the tuning parameters set to $0.5$, the lasso SIR estimator selected seven genes, the two-stage lasso selected nine genes, and the proposed two-stage lasso SIR estimator selected six genes. Only one gene was chosen with a very high selection probability across all the three methods, while one additional gene was commonly selected by both the two-stage lasso SIR and two-stage lasso estimators.
@@
Figure 1: Stability selection paths for three different methods applied to the mouse obesity data based on 100 random subsamples. With the maximum selection probability of at least 0.5, the gene selected by all the three methods is displayed in red, while the additional gene only selected by both two-stage methods is displayed in blue.
@@ 7.2 Measurement error: NHANES data
The National Health and Nutrition Examination Survey (NHANES) aims to track the health and nutritional status of individuals in the United States over time (Centers for Disease Control and Prevention,, 2018). During the 2009-2010 survey period, participants were interviewed and asked to provide their demographic background, nutritional habits, and to undertake a series of health examinations. To assess the nutritional habits of participants, dietary data were collected using two 24-hour recall interviews, wherein the participants self-reported the consumed amounts of a set of food items during the 24 hours before each interview. Based on these recalls, daily aggregated consumption of water, food energy, and other nutrition components such as total fat and total sugar consumption were computed. In this section, we revisit the data application in Nghiem et al., 2023a , seeking to understand the relationship between total cholesterol level ($y$) for $n=3343$ women and their long-term intakes of $p=42$ nutrition components, particularly a linear combination/s of these components carrying the full regression information, and whether these combinations are sparse in nature.
@@
Let $\mathbf{W}_{ij}$ denote the vector of 42 nutritional covariates at the $j$th interview for $j=1,2$. Dietary recall data are well-known to be subject to measurement errors, so we assume $\mathbf{W}_{ij}=\mathbf{X}_{i}+{E}_{ij}$, where $\mathbf{X}_{i}$ represents the $i$th latent long-term nutritional intake, and ${E}_{ij}$ denotes an independent additive measurement error term following a multivariate Gaussian distribution with zero mean and covariance $\bm{\Sigma}_{u}$. We used an instrumental variable approach to account for this measurement error. Specifically, following Carroll et al., (Section 6.5.1, 2006), we treated the first replicate $\mathbf{W}_{i1}=\mathbf{W}_{i}$ as a surrogate for $\mathbf{X}_{i}$ and the second replicate $\mathbf{W}_{i2}=\mathbf{Z}_{i}$ as an instrument. Note in this case, the number of instruments satisfies $q=p$. We then applied the linear instrumental variable model $\mathbf{X}_{i}=\bm{\Gamma}^{\top}\mathbf{Z}_{i}+\mathbf{U}_{i}$, where $\bm{\Gamma}$ is allowed to be a $q\times p$ column-sparse matrix. It follows that we can write $\mathbf{W}_{i}=\bm{\Gamma}^{\top}\mathbf{Z}_{i}+\tilde{\mathbf{U}}_{i}$ where $\tilde{\mathbf{U}}_{i}=\mathbf{U}_{i}+\mathbf{E}_{i1}$. We then applied the proposed two-stage lasso SIR estimator replacing $\mathcal{X}$ with $\mathcal{W}$. All three methods for selecting the structural dimension described in Section 6 selected $\hat{d}=1$; this was also consistent with the analysis in Nghiem et al., 2023a .
@@
We first examined the estimated matrix coefficients $\widehat{\bm{\Gamma}}$ from the first stage of the two-stage lasso SIR estimation procedure. Figure 4 in Appendix D shows that the largest element in each column of $\widehat{\bm{\Gamma}}$ is typically the diagonal element, so the most important instrument to model the consumption of any dietary component in the second interview is, unsurprisingly, the consumption of the same component in the first interview. Overall, each column of $\widehat{\bm{\Gamma}}$ is relatively sparse, with the number of non-zero coefficients in each column ranging from 3 to 20. Next, we applied the same stability selection approach as in Section 7.1 to the proposed two-stage lasso SIR estimator, and compared the resulting paths with the lasso SIR and two stage lasso estimators. With the maximum selection probability across all the tuning parameters being at least 0.5, the lasso SIR estimator selected five nutritional covariates, the two-stage lasso estimator selected five nutritional covariates, while the proposed two-stage lasso SIR estimator selected six nutritional covariates. Figure 2 presents the selection path as a function of the tuning parameter $\mu_{2}$ for each nutrition component. The two components commonly selected by all the three estimators were vitamin D and vitamin K, while both the two-stage lasso SIR and two stage lasso estimators additionally select one more common component in folic acid. Vitamin D is the component with the largest estimated coefficients in both the two-stage lasso SIR and two-stage lasso estimators, and was the second largest in the lasso SIR estimator.
@@ 8 Discussion
In this paper, we propose a two-stage lasso SIR estimator for simultaneous sufficient dimension reduction and variable selection in the presence of endogeneity. We establish estimation and selection consistency of the proposed estimator, allowing the dimensionality of both the covariates and instruments to grow exponentially with sample size. Numerical studies demonstrate the superior performance of the proposed estimator compared with existing methods that ignore non-linearity and/or endogeneity in the outcome model.
Future research can extend the proposed method to accommodate a possible non-linear relationship between the covariates and the instruments in the first stage. Different sparsity structures on the coefficient matrix $\bm{\Gamma}$, such as a low-rank matrix, can also be considered (reinsel2022multivariate). We can also examine endogeneity in other sufficient dimension reduction methods, such as inverse regression models (Bura et al.,, 2022), and finally, we can adapt the methods to more complex settings involving endogeneity, such as SDR for clustered or longitudinal data (hui2022sufficient; Nghiem and Hui,, 2024).
Appendix A Performance of Lasso SIR under endogenity
@@
[…]
@@
[…]
@@
For each generated sample, we computed the Lasso SIR estimator $\hat{\bm{\beta}}^{*}$ with the tuning parameter $\mu$ selected by 10-fold cross-validation, as implemented by the LassoSIR package. We reported the estimation error in estimating the column space spanned by $\bm{\beta}$, as measured by $\lVert\mathcal{P}(\bm{\beta})-\mathcal{P}(\hat{\bm{\beta}})\rVert_{F}$. For variable selection performance, we computed the area under the receiver operating curve (AUC); an AUC of one indicates a clear separation between the estimates of non-important variables and the those of important variables in $\hat{\bm{\beta}}^{*}$.
Table 4 shows that the performance of the Lasso SIR estimator aligns with our theoretical results. In the first case, the correlation vector $\bm{\Sigma}_{\mathbf{X},\varepsilon}$ is proportional to the true $\bm{\beta}$, so the Lasso SIR maintains both estimation and variable selection consistency, as evidenced by a decreasing estimation error when $n$ increases and an AUC of 1. In the second case, the Lasso SIR estimator gives a consistent estimate of the space spanned by $\bm{\eta}^{*}=(1.5,0.5,0,0)/\sqrt{2.25}$, which has a different direction but the same support compared to the true $\bm{\beta}$. Hence, the Lasso SIR is not estimation consistent, but still has an AUC of 1. In the last case, the Lasso SIR estimator gives a consistent estimate for the space spanned by $\bm{\eta}^{*}=(0.5,0.5,0.5,0.5)^{\top}$, which has both a different direction and a different support from those of the true $\bm{\beta}$. As a result, the Lasso SIR estimator is not selection or estimation consistent.
Table 4: Performance of the Lasso SIR estimator under endogeneity, as measured by the estimation error $\lVert\mathcal{P}(\bm{\beta})-\mathcal{P}(\hat{\bm{\beta}}^{*})\rVert_{F}$ and variable selection performance (AUC).
@@
Appendix B Technical Proofs
B.1 Proof of Proposition 1
Without loss of generality, assume the expectation of $\text{E}(\mathbf{X},\varepsilon)=0$, the marginal covariance matrix $\bm{\Sigma}_{X}=\mathbf{I}_{p}$, and the marginal variance $\operatorname{\text{Var}}(\varepsilon)=1$. In this case, it follows that
[…]
@@ B.2 Proof of Proposition 2
$\frac{1}{n}\lVert\tilde{\bm{y}}^{*}_{k}-\mathcal{X}\tilde{\bm{\beta}}^{*}_{k}-\mathcal{X}\bm{\delta}_{k}^{*}\rVert^{2}+\mu_{k}\|\tilde{\bm{\beta}}^{*}_{k}+\bm{\delta}_{k}^{*}\|_{1}\leq\frac{1}{n}\lVert\tilde{\bm{y}}^{*}_{k}-\mathcal{X}{\bm{\beta}}^{*}_{k}\rVert^{2}+\mu_{k}\|\tilde{\bm{\beta}}^{*}_{k}\|_{1}.$
@@
By construction, we have $n^{-1}\mathcal{X}^{\top}\tilde{\mathbf{y}}_{k}^{*}=\hat{\bm{\eta}}_{k}^{*}$, the eigenvector associated with the $k$th largest eigenvalue of $\hat{\bm{\Lambda}}^{*}$. By the law of large numbers, we have $\hat{\bm{\Lambda}}^{*}\xrightarrow{p}\bm{\Lambda}^{*}$, so $\lVert\hat{\bm{\eta}}_{k}^{*}-\tilde{\bm{\eta}}_{k}^{*}\rVert=o_{p}(1)$. Furthermore, as $n\to\infty$, we have $n^{-1}{\mathcal{X}^{\top}\mathcal{X}}\xrightarrow{p}\bm{\Sigma}_{\mathbf{X}}$ so $n^{-1}\mathcal{X}^{\top}\mathcal{X}\tilde{\bm{\beta}}_{k}^{*}\xrightarrow{p}\tilde{\bm{\eta}}_{k}^{*}$. Therefore, as $n\to\infty$. the vector $\bm{\delta}_{k}^{*}$ satisfies
@@
Since $\mu_{k}=o(1)$ and $\lVert\hat{\bm{\eta}}_{k}^{*}-\tilde{\bm{\eta}}_{k}^{*}\rVert=o_{p}(1)$, we have $\lVert\bm{\delta}_{k}^{*}\rVert=o_{p}(1)$, so $\mathcal{P}(\mathbf{B}_{k}^{*})\xrightarrow{p}\mathcal{P}(\widetilde{\mathbf{B}}^{*})=\mathcal{P}({\mathbf{B}}^{*})$ as $n\to\infty$.
B.3 Proof of Lemma 1
The instrument $\mathbf{Z}$ is independent of both $\varepsilon$ and $\mathbf{U}$, so the conditional expectation $\widehat{\mathbf{X}}=E(\mathbf{X}\mid\mathbf{Z})=\bm{\Gamma}\mathbf{Z}$ is also independent of both $\varepsilon$ and $\mathbf{U}$. By the properties of conditional expectation, we have